simplify
ln(x^3-x^2)

Found 3 solutions by HyperBrain, MissionPossible, ikleyn:
Answer by HyperBrain(694)

ln(x3-x2)=ln(x(x+1)(x-1))=ln x +ln (x+1) +ln (x-1)

Power up,
HyperBrain!

Stop this mess! Visit these sites: http://uselessplastic.wordpress.com
http://uselessplastic.multiply.com

Answer by MissionPossible(18)

ln(x^3-x^2) = ln(x^2(x-1)) = ln (x^2) + ln (x-1) = 2 ln x + ln (x-1)

Answer by ikleyn(53937)

.
simplify ln(x^3-x^2).
~~~~~~~~~~~~~~~~~~~~~~~~~


The formula deriving and the answer in the post by @HyperBrain both are incorrect.

The correct solution is THIS

ln(x^3-x^2) = ln(x^2*(x-1)) = 2ln(x) + ln(x-1)   for   x > 1.         ANSWER